If 10.0 g of substance A decomposes for 100 min, the mass of A remaining unreacted is found to be 1.25g. What is the half life of this reaction if it follows first order kinetics?
Answer #1

A -------> product

At the beginning of the reaction , the mass of A is 'a' gm . At time 't' the amount of A that has been decomposed is 'x' gm.

From the first order reaction , we can write

K = \frac{1}{t}ln\left ( \frac{a}{a-x} \right ) .,..,.............. (1)

Here K = rate constant

Here given that

a = initial mass = 10.0 g

t =time = 100 min

(a-x) = unreacted mass at time t = 1.25 g

Putting this value  

K = \frac{1}{100min}ln\left ( \frac{10g}{1.25g} \right )

K = 0.0208 min -1

For half life , we know the the decomposition of half mass .

Hence x= ( a/2) when t = t​​​​​​1/2

Putting this value in equation (1) , we get

K = \frac{1}{t_{1/2}}ln\left ( \frac{a}{a-a/2} \right )

K = \frac{1}{t_{1/2}}ln\left ( \frac{a}{a/2} \right )

K = \frac{1}{t_{1/2}}ln2

{t_{1/2}}=\frac{1}{K} ln2

Putting the k value

{t_{1/2}}=\frac{1}{0.0208 min^{-1}} ln2

{t_{1/2}}= 33.32 min

Hence half of life of this reaction is 33.32 min .